OpenStudy (anonymous):

Will Give medals! Which function below has the smallest slope? A Graph with a line going through -5 on the Y Axis and Postive 2 and a half on the X axis. f(x) f(x) = 2x - 3 h(x) x h(x) -1 3 1 7 3 11

OpenStudy (anonymous):

could you please type the question properly its hard to understand

OpenStudy (anonymous):

There are three solutions I am looking for the smallest slope. Answer 1 Is A Graph with a line going through -5 on the Y Axis and Positive 2 and a half on the X axis. Answer 2 f(x) f(x) = 2x - 3 answer 3 h(x) x h(x) -1 3 1 7 3 11

OpenStudy (anonymous):

f(x) = 2x - 3 is of the form y= mx+b where m is the slope so slope of f(x) = 2

OpenStudy (anonymous):

if a line is going through -5 on the y axis and +2 on the x axis then (0,-5) and (2,0) are two points that exist on the line the slope m = (y2-y1)/(x2-x1) = 5/2 = 2.5

OpenStudy (anonymous):

i cant understand what answer three is